A screw jack, 25% efficient and having a screw of pitch 0.4cm is used to raise a load through a certain height. If I’m the process the handle turns through a circle of radius 40cm, calculate the

i) velocity ratio of the machine

ii) The mechanical advantage of the machine

iii) Effort required to rise a load of 1000N with the machine. (take π = 3.14)


1
Expert's answer
2021-11-14T17:11:38-0500

i) velocity ratio of the machine:


"VR=\frac{2\pi r}{p}=\frac{2\pi·40}{0.4}=628."

ii) The mechanical advantage of the machine:


"MA=\frac{2\pi r}{p}=\frac{2\pi·40}{0.4}=628."

iii) Effort required to rise a load of 1000N with the machine:


"E=\frac{F}{MA·\eta}=\frac{1000}{628·0.25}=6.4\text{ N}."